OpenStudy (anonymous):

Solve 2cos^2Θ/2 = cos2Θ for 0º≤Θ≤360º.

OpenStudy (anonymous):

I tried graphing it to solve, but I must be doing something wrong because none of my answers match

OpenStudy (jdoe0001):

\(\bf \cfrac{2cos^2(2\theta)}{2}=cos(2\theta)\quad ?\)